# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ 5.2. Identification of Pertinent Story Elements
For each scenario and each potential value of the elements of interest (i.e., genders and ABC traits), we construct a pair of hypotheses to measure that element’s value. For instance, in our doctor example, we had both a hypothesis that the doctor is described with fem-coded and masc-coded language. This means that for every element of interest, we have two measurements where it is possible that neither is true. We find that the model marks both options as true in only $0.2\%$ (eight total) of these annotations and correct them manually. In the case of gender, we do not annotate for identities beyond the binary cases. We expected that participants would not make characters explicitly transgender or non-binary, and we could not find any such cases through a manual evaluation of a small sample of stories.
@@
To prompt the model for story and word-level annotations, we provide the model with simple instructions, in-context examples to demonstrate the task (Figure 17), and the hypothesis and a full or partial story. The partial stories are used to collect word-by-word measurements of the elements of interest at every [^8] Due to technical limitations, server lag caused fast typing participants to, at times, not receive an updated set of suggestions for every word they inputted leading to some nonsensical model-suggested word continuations. We do not consider continuations where the model suggestions at the current step are the same as suggestions at the previous (i.e., when the server has lagged). step, both for the words that are included in the final story and for the model-suggested words that are rejected. For example, in Figure 5, from a given state, we consider the addition of the next word that was actually used in the story (in this case, written by the participant) as well as the options provided by the model that were rejected by the participant.
@@
[…]
Example word-level annotations. We start with the prefix “Joe was happy to be marrying” which the model marks as having a very low probability of Joe’s partner being a man (or masc-coded). We then consider the addition of the next word included in the story (“Sarah”) and the next words suggested by the model (“his”, “, “Steve”, and “the”). We collect similar model predictions with each of these words added. We see that adding in “Sarah”, “his”, or “the” keeps the probability of the partner being a man low while “Steve” flips this completely. This shows that the word “Steve” specifies the partner’s gender.
@@
Based on a manual examination of the data, we empirically choose $0.8$ as the cutoff point for determining whether an attribute is present in the story. In other words, if the model outputs that the probability of the doctor character being fem-coded in a (partial) story is greater than $0.8$, then we consider the (partial) story as having a fem-coded doctor in it. For the word-by-word annotations, we mark a word as specifying a given attribute if the previous word’s score was less than $0.8$, the new word’s score is greater than $0.8$ and the difference between them is greater than $0.3$. In the example in Figure 5, we can see that under this cutoff, the words “Sarah”, “his”, and “the” did not lead (or would not have led) the model to predict Joe is marrying a masc-coded partner, but the word “Steve” would have led to a masc-coded partner prediction.
@@
We employ some manual and automated cleaning to these annotations, as the purpose here is not to evaluate the LLM’s ability to annotate these (parts of) stories but to obtain a reliable set of annotations of gender and ABC traits. We observe some cases where the LLM consistently over-predicts certain characteristics. For example, in the Wedding scenario, the partial story “Sherry was happy to be marrying her”, the LLM understandably produces that Sherry’s partner is a woman as though “her” is an object pronoun (i.e., the pronoun refers to Sherry’s partner). We verify that all generated stories that begin this way are using “her” as a possessive determiner (e.g., as in “her fiancee”) and remove all such cases from the set of words that would determine that Sherry’s partner is fem-coded.
@@
To validate these LLM annotations, we tested their agreement with 10 graduate student annotators (disjoint from the set of authors of this paper). They were asked to annotate $560$ story-attribute pairs total covering every axis of interest and were paid \$$\$5$ for an median compensation rate of \$$\$15.79$ per hour. For more complete details including instructions, see subsection C.3. Pooling annotations between human annotators, we find an overall agreement level per Cohen’s Kappa of $\kappa_{\text{all}}=0.768$ which constitutes “substantial agreement” (Landis and Koch, 1977) between humans and the LLM annotator. For the gender annotations, we find an agreement of $\kappa_{\text{gender}}=0.782$. For the other ABC traits (likeability, assertiveness, etc), the annotations are only slightly more subjective, where we find an agreement of $\kappa_{\text{ABC}}=0.757$.
@@ 6. Results
We then discuss the influences of biased predictive text in three writing scenarios in more detail at the story and word level for a subset of the scenarios from Table 1: Detective (subsection 6.2, Wedding (subsection 6.3), and President (subsection 6.4). We selected these three scenarios for the body of the paper to cover a variety of gender configurations. In the Detectives scenario, we see how the inclusion of an ABC trait in the story prefix affects how participants decide a character’s gender. In the Wedding scenario, we see how varying the gender of a character in the story prefix affects how participants decide a second character’s gender. And in the President we consider how participants decide a character’s gender and ABC trait jointly. Similar analyses of the remaining scenarios is given in Appendix A and follows similar trends, though in some cases with more mixed conclusions.
@@
Unless otherwise stated, all comparisons in this analysis were made using independent t-tests. We perform Benjamini-Hochberg correction to avoid multiple testing and report the adjusted p-values with $p_{\text{FDR}}$. Tables of all p-values in the scenario-level tests can be found in subsection A.5. All error bars in our figures show $90\%$ confidence intervals.
@@ 6.1. Summary of Gender Effects at the Story Level
We summarize our results surrounding gender alone in overall stories for scenarios where the participant (not the story prefix) can decide a character’s gender Table 2. In all scenarios but “Joe’s” wedding, suggestions of a fem-coded character are anti-stereotypical. Here, we see that masc-coded suggestions never significantly change the rates of writing characters of any gender when compared to stories written without suggestions. On the other hand, fem-coded suggestions significantly (or marginally) decrease the rate of masc-coded characters or increase the rate of fem-coded characters. However, even with this change, we still see at least marginally more fem-coded characters than masc-coded characters in every scenario under every setting, including fem-coded suggestions.
In the scenario of “Joe’s” wedding, suggestions of a masc-coded character are anti-stereotypical. Here we see that masc-coded suggestions do not significantly affect the rates of writing characters of any gender when compared to stories written without suggestions, but we see an insignificant trend towards writing fewer fem-coded partners with masc-coded suggestions (as discussed more in subsection 6.3). Regardless of the presence or type of suggestions, we always see significantly more fem-coded partners for “Joe” than masc-coded.
@@
(a) Comparison of rates of writing characters pro-stereotypical (or anti-stereotypical gender) with pro-stereotypical or anti-stereotypical suggestions and no suggestions as a baseline. Changes marked with an arrow are statically significant.
@@
Detective 006.44×6.$4$ 10.29×10.29$\times$10.29 × 001.37×1.37$\times$1.37 ×
President 14.67×14.67$\times$14.67 × $3$ 002.38×2.38$\times$2.38 ×
Doctor 006.33×6.$3$ 007.56×7.56$\times$7.56 × 001.23×1.23$\times$1.23 ×
Wedding (Joe) 12.00×12.$0$ 11.17×11.17$\times$11.17 × $0$
Wedding (Sherry) 005.67×5.67$\times$5.67 × 008.25×8.25$\times$8.25 × 001.44×1.$4$
(b) How many times more pro-stereotypically gendered characters were written than anti-stereotypically gendered characters with various suggestions. Numbers shown in gray are statistically significant. No number is $<1$, meaning all stories had at least as many pro-stereotypically gendered characters than anti-stereotypically gendered characters, regardless of condition.
@@
Because we see a lesser rate of acceptance of anti-stereotypical suggestions in our writing scenarios, this means that a “debiased” predictive text system that suggests fem-coded and masc-coded characters at a equal rate will not yield an equal proportion of stories with fem-coded and masc-coded characters. In Figure 6, we show how as the rate of theoretical users shown anti-stereotypical suggestions increases, the expected proportion of anti-stereotypically gendered characters also increases. [^9] In our study design, we also only consider cases where the LLM is explicitly prompted to suggest one character attribute or another. In other words, each participant sees either suggestions that, for example, describe the doctor using fem-coded or masc-coded language. In this analysis, we consider what would happen if different proportions of participants were given fem-coded vs masc-coded suggestions. Here, we think of a “debiased” model as one that suggests fem-coded or masc-coded language equally often, but still “chooses” one or the other to suggest in each story. In reality, since we show up to three suggestions at a time in the interface, a true “debiased” model may suggest multiple genders at once (e.g., having both “she” and “he” among the top three suggestions). Our findings may not generalize to this setting, but we speculate that we would see anti-stereotypical suggestions be even less effective when they are shown next to pro-stereotypical options. We can see that there is no scenario where any configuration of suggestions yields gender parity. However, we do see that even a small proportion of anti-stereotypical suggestions is enough to increase the rate of anti-stereotypically gendered characters over the observed rate for stories written without suggestions. For example, in the Doctor scenario, raising the rate of fem-coded suggestions over $6.4\%$ fem-coded suggestions would yield, in expectation, more fem-coded doctors than writing without suggestions.
@@
Figure 6. Expected rates of anti-stereotypically gendered characters in human-written stories (y-axis) with no suggestions (left) and as the proportion of anti-stereotypical predictive text suggestions increases (x-axis, right). Note that everything falls below $y=0.5$ or “perfect” gender parity. In our study, we measure the two extremes ($x=0$ and $x=1$) for each scenario; the other points are interpolated. In general, a model that suggests entirely pro-stereotypical text yields stories that are only slightly (if at all) more stereotyped than with no suggestions. And a model that suggests entirely anti-stereotypical text increases the rate of anti-stereotypical stories, but never so much so as to even reach parity with pro-stereotypical stories. Note that the variance in $x$ value on the left plot is for visual clarity only.
[…]
Line graphs showing how the expected rates of anti-stereotypically gendered characters as the proportion of anti-stereotypical suggestions increases. We show results for all scenarios where the participant can specify a character’s gender (detectives, president, doctor, and wedding). In all cases, we see that even with $100\%$ anti-stereotypical suggestions, we still expect to yield less than $50\%$ anti-stereotypical characters. We also include the proportion of anti-stereotypically gendered characters we observed in stories written without suggestions. We can see that yielding more anti-stereotypically gendered characters than the “no suggestions” condition requires a relatively small proportion of anti-stereotypcial suggestions: $13.8\%$ for detectives, $15.5\%$ for presidents, $6.4\%$ for doctors, and $14.4\%$ for weddings with “Sherry”. For weddings with “Joe”, we surprisingly see slightly more anti-stereotypically gendered characters (i.e., masc-coded partners) even when only providing pro-stereotypical suggestions.
@@ 6.2. Scenario: Detectives
In this scenario, participants continue from a story prefix that describes Detective Wilson’s partner as either trustworthy or untrustworthy. The model then suggests that Detective Wilson’s partner is fem-coded or masc-coded (see examples in Table 5). According to Cao et al. (2022b), American annotators view men as comparatively less trustworthy than women. We treat cases with masc-coded detectives to be pro-stereotypical and cases with trustworthy fem-coded detectives (and untrustworthy masc-coded detectives) to be more pro-stereotypical than their untrustworthy fem-coded (and trustworthy masc-coded) counterparts.
@@ 6.2.1. Effects on Gender Alone
First, we consider the effects of suggestion on gender, regardless of the trustworthiness of Detective Wilson’s partner. We first analyze gender at the story-level, then the relationship between gender and word-level reliance.
@@
[…]
Story-level gender breakdown of detective partners. Considering all stories, we see that the gender breakdown with no suggestions and masc-coded suggestions are very similar (both with much more masc-coded detectives than fem-coded). With fem-coded suggestions, we see more fem-coded detectives, but there are still more masc-coded detectives than fem-coded. When we break this down further by the trustworthiness of the partner detective, we see more masc-coded detectives when they are untrustworthy.
At the Story Level
First, we compare the proportion of stories with partners of a given gender that were written with vs without suggestions (Figure 7). Here, we see no significant differences when comparing gender rates without suggestions to rates with masc-coded suggestions (masc-coded partners: $t(242)=0.899$, $p_{\text{FDR}}\approx 0.5681$; fem-coded partners: $t(242)=-1.012$, $p_{\text{FDR}}\approx 0.5254$). When the model suggests that the partner should be fem-coded, we see significantly fewer masc-coded partners ($t(234)=-4.0$, $p_{\text{FDR}}\approx 0.0007$) and significantly more fem-coded partners ($t(234)=4.191$, $p_{\text{FDR}}\approx 0.0003$). However, even with these changes, we still see significantly more masc-coded partners than fem-coded in all conditions (no suggestions: $t(144)=-10.934$, $p_{\text{FDR}}<0.0001$; masc-coded suggestions: $t(340)=-21.727$, $p_{\text{FDR}}<0.0001$; fem-coded suggestions: $t(324)=-2.575$, $p_{\text{FDR}}\approx 0.0408$).
@@
We further assess participants’ reliance on model suggestions at the word level (Table 3). When considering all words suggested by the model in the fem-coded and masc-coded conditions, we see that participants are significantly more likely to write new words or edit model suggestions in the fem-coded detective condition (H2; $t(7091)=4.724$, $p_{\text{FDR}}<0.0001$), but we see no significant trend when we constrain this to only the story words that determine the second detective’s inferred gender (H2a; $t(302)=-0.069$, $p_{\text{FDR}}\approx 0.9634$). We also see that participants are significantly less likely to accept model suggestions that would make the second detective fem-coded (H2b; $t(834)=6.729$, $p_{\text{FDR}}<0.0001$).

### v2

@@ 5.2. Identification and Measurement of Pertinent Story Elements
For each scenario and each potential value of the elements of interest (i.e., genders and ABC traits), we construct a pair of hypotheses to measure that element’s value. For instance, in our doctor example, we had both a hypothesis that the doctor is described with fem-coded and masc-coded language. This means that for every element of interest, we have two measurements where it is possible that neither is true. We find that the model marks both options as true in only $0.2\%$ (eight total) of these annotations and correct them manually. We do not annotate for gender identities beyond binary ones. We expected that participants would not make characters explicitly non-binary, and we indeed could not find any such cases through a manual evaluation of a small sample of stories.
@@
To prompt the model for story and word-level annotations, we provide the model with simple instructions, in-context examples to demonstrate the task, the hypothesis, and the full or partial story premise (See Figure 19). The partial stories are used to collect word-by-word measurements of the elements of interest at every [^9] Due to technical limitations, server lag caused fast typing participants to, at times, not receive an updated set of suggestions for every word they inputted leading to some nonsensical model-suggested word continuations. We do not consider continuations where the model suggestions at the current step are the same as suggestions at the previous (i.e., when the server has lagged). step, both for the words that are included in the final story and for the model-suggested words that are rejected. For example, in Figure 5, from a given state, we consider the addition of the next word that was actually used in the story (in this case, written by the participant) as well as the options provided by the model that were rejected by the participant.
@@
Example word-level annotations starting from the prefix “Joe was happy to be marrying.” The figure shows model probabilities for the partner being masculine-coded after adding the next word “Sarah” or model suggestions “his,” “Steve,” and “the.” Adding “Sarah,” “his,” or “the” keeps the probability of the partner being a man low, while adding “Steve” increases the probability, indicating a masculine-coded partner.
@@
Based on a manual examination of the data, we empirically choose $0.8$ as the probability cutoff point for determining whether an attribute is present in the story. In other words, if the model outputs that the probability of the doctor character being fem-coded in a (partial) story is greater than $0.8$, then we consider the (partial) story as having a fem-coded doctor in it. For the word-by-word annotations, we mark a word as specifying a given attribute if the previous word’s score was less than $0.8$, the new word’s score is greater than $0.8$ and the difference between them is greater than $0.3$. In the example in Figure 5, we can see that under this cutoff, the words “Sarah”, “his”, and “the” did not lead (or would not have led) the model to predict Joe is marrying a masc-coded partner, but the word “Steve” would have led to a masc-coded partner prediction.
@@
We employ manual and automated cleaning on these annotations, as the purpose is not to evaluate the LLM’s ability to annotate these (parts of) stories but to obtain a reliable set of annotations of gender and ABC traits. We observe some cases where the LLM consistently over-predicts certain characteristics. For example, in the Wedding scenario, the partial story “Sherry was happy to be marrying her”, the LLM understandably predicts that Sherry’s partner is a woman as though “her” is an object pronoun (i.e., the pronoun refers to Sherry’s partner). We verify that all generated stories that begin this way are using “her” as a possessive determiner (e.g., as in “her fiancee”) and remove all such cases from the set of words that would determine that Sherry’s partner is fem-coded. For more details about how model annotations were generated, see Appendix D.
@@
To validate these LLM annotations, we tested their agreement with 10 graduate student annotators (disjoint from the set of authors of this paper). They were asked to annotate $560$ story-attribute pairs total covering every axis of interest and were paid $\mathdollar 5$ for an median compensation rate of $\mathdollar 15.79$ per hour. For more complete details including instructions, see subsection C.3. Pooling annotations between human annotators, we find an overall agreement level per Cohen’s Kappa of $\kappa_{\text{all}}=0.768$ which constitutes “substantial agreement” (Landis and Koch, 1977) between humans and the LLM annotator. For the gender annotations, we find an agreement of $\kappa_{\text{gender}}=0.782$. For the other ABC traits (likability, assertiveness, etc), we find a slightly lower agreement of $\kappa_{\text{ABC}}=0.757$ perhaps as these traits are more subjective than gender.
These story-level and word-level annotations are then used as outcome measurements in H1, H2a, and H2b. In H1, we consider how often specific stereotype-relevant content is present in overall stories. For example, we consider how likely a president character is to be described using fem-coded language, depending on the presence or type of suggestions. In H2a, we consider how often the words in the final story (especially those that mark stereotype-relevant features) were suggested by the model. For example, we measure how often any word in the final story was accepted exactly from a model suggestion when the model is prompted to suggest the president uses fem-coded language. We consider the same measurement on the subset of words that mark gender (where this subset is chosen based on the change in LLM annotator confidence in the president’s gender). Finally, in H2b, we consider when the model suggests words that mark stereotype-relevant features, how often the participants are to accept them. For example, we measure how often participants accept model suggested words when these words would mark a the president as fem-coded. These story-level and word-level annotations are then used as outcome measures in the analysis below. Together, these annotations provide the outcome measures used in the analyses that follow.
@@ 6. Results
We then discuss the influences of biased predictive text in more detail at the story and word level for a subset of three scenarios from Table 1: Detective (subsection 6.2), Wedding (subsection 6.3), and President (subsection 6.4). We selected these three scenarios for the body of the paper to cover a variety of trait configurations. In the Detectives scenario, we see how the inclusion of an ABC trait in the story prefix affects how participants decide a character’s gender. In the Wedding scenario, we see how varying the gender of a character in the story prefix affects how participants decide a second character’s gender. And in the President we see how participants decide a character’s gender and ABC trait jointly. Similar analyses of the remaining scenarios are in Appendix A and follows similar trends, though in some cases with more mixed conclusions.
@@
Unless otherwise stated, all comparisons in this analysis were conducted using independent t-tests, with effect sizes reported as Cohen’s $d$. We perform Benjamini-Hochberg correction and report the adjusted p-values with $p_{\text{FDR}}$. Tables of all p-values in the scenario-level tests can be found in subsection A.5. All error bars in our figures show $90\%$ confidence intervals.
@@ 6.1. Summary of Gender Effects at the Story Level
(a) Comparison of rates of writing characters pro-stereotypical (or anti-stereotypical gender) with pro-stereotypical or anti-stereotypical suggestions and no suggestions as a baseline. Changes marked with an arrow are statically significant. In our scenarios, it is possible for stories to include neither the pro-stereotypical nor the anti-stereotypical trait (e.g., the doctor’s gender is never specified). This means that a change to the rate of pro- or anti-stereotypical stories does not necessitate a corresponding change to the other.
@@
[…]
(b) How many times more pro-stereotypically gendered characters were written than anti-stereotypically gendered characters with various suggestions. Numbers shown in gray are statistically significant. No number is <1<1, meaning all stories had at least as many pro-stereotypically gendered characters than anti-stereotypically gendered characters, regardless of condition.
@@
Figure 6. Expected rates of anti-stereotypically gendered characters in human-written stories (y-axis) with no suggestions (left) and as the proportion of anti-stereotypical predictive text suggestions increases (x-axis, right). We can see that even with exclusively anti-stereotypical suggestions, we predict that gender parity falls below $y=0.5$ or “perfect” gender parity. In our study, we measure the two extremes ($x=0$ and $x=1$) for each scenario and calculate the fraction of stories with characters of each gender. The predicted values from other distributions of suggestions are drawn from a linear interpolation between those points. In general, a model that suggests entirely pro-stereotypical text ($x=0$) yields stories that are only slightly (if at all) more stereotyped than with no suggestions. And a model that suggests entirely anti-stereotypical text ($x=1$) increases the rate of anti-stereotypical stories, but never so much so as to even reach parity with pro-stereotypical stories. Note that the variance in $x$ value on the left plot is for visual clarity only.
Line graphs showing expected rates of anti-stereotypically gendered characters as the proportion of anti-stereotypical suggestions increases for detectives, presidents, doctors, and weddings. Each curve shows the observed rates with no suggestions as a reference. Approximate thresholds where curves surpass the no-suggestion baseline are indicated:  14% for detectives,  15% for presidents,  6% for doctors, and  14% for weddings with “Sherry.” The curve for weddings with “Joe” shows a slight increase in masculine-coded partners even with pro-stereotypical suggestions.
We summarize our results on gender in overall stories for scenarios where participants, rather than the story prefix, determined a character’s gender in Table 2. In all scenarios except “Joe’s” wedding, fem-coded character suggestions are anti-stereotypical. We find that pro-stereotypical suggestions have no significant effects when compared to writing without suggestions (2(a), left). By contrast, anti-stereotypical suggestions significantly or marginally shift writing toward anti-stereotypical characterization or away from pro-stereotypical characterization (2(a), right). Still, in every scenario, pro-stereotypical characters remain (often significantly) more common than anti-stereotypical ones, despite exclusively anti-stereotypical suggestions (2(b)).
These experiments test extreme cases: predictive text suggestions that are entirely pro- or anti-stereotypical. As discussed in section 3, there is no single definition of “fair” predictive text suggestions or distributions of stories. Importantly, our anti-stereotypical condition should be understood as an upper bound, stronger than what most researchers or practitioners would consider a “fair” or “debiased” model. [^10] In our study design, we also only consider cases where the LLM is explicitly prompted to suggest one character attribute or another. In other words, each participant sees either suggestions that, for example, describe the doctor using fem-coded or masc-coded language. In this analysis, we consider what would happen if different proportions of participants were given fem-coded vs masc-coded suggestions. Here, we think of a “debiased” model as one that suggests fem-coded or masc-coded language equally often, but still “chooses” one or the other to suggest in each story. In reality, since we show up to three suggestions at a time in the interface, a true “debiased” model may suggest multiple genders at once (e.g., having both “she” and “he” among the top three suggestions). Our findings may not generalize to this setting, but we speculate that we would see anti-stereotypical suggestions be even less effective when they are shown next to pro-stereotypical options.
To examine more realistic bias configurations, Figure 6 shows how the expected proportion of anti-stereotypical stories changes for less extreme proportions of anti-stereotypical suggestions. We find no setting that would yield perfect parity in depicting fem-coded and masc-coded characters, even when suggestions are exclusively anti-stereotypical. However, even small amounts of anti-stereotypical suggestions are enough to increase anti-stereotypical writing above the baseline (unassisted) rate. For instance, in the Doctor scenario, $6.4\%$ fem-coded suggestions are estimated to be sufficient to increase fem-coded doctors over unassisted writing.
In cases where developers are targeting a specific distribution of stories (e.g., one that matches some real-world gender distribution), we see that it may be possible to yield an intended distribution; however, this would require the use of a model that is more “anti-stereotypical” than the developer is targeting. For example, in the United States, $26.3\%$ of detectives were women in 2023.111111https://www.bls.gov/cps/cpsaat11.htm A model that suggests $26.3\%$ fem-coded detectives would yield only an expected $17.7\%$ fem-coded detectives in co-written stories. To reach $26.3\%$ fem-coded detectives, we expect that model would need to suggest $52.2\%$ fem-coded detectives.
Regardless of whether a developer conceives of a “fair” model as one satisfying parity constraints or matching some real-world demographic distribution, we consistently find that the difference in human acceptance of pro-stereotypical and anti-stereotypical model suggestions will lead to a final gender distribution of human-AI stories where anti-stereotypical stories do not significantly outnumber pro-stereotypical stories (and in fact, are often still significantly outnumbered by pro-stereotypical stories). To achieve a “fair” gender distribution in human-AI co-written stories, developers would need to suggest anti-stereotypical completions more often than they think is “fair”, and depending on their definition of a “fair” distribution, the desired outcome may not be possible through predictive text suggestions alone. Thus, while increasing the rate of anti-stereotypical suggestions can help encourage users to compose more anti-stereotypical writing, interventions that focus exclusively on debiasing model suggestions may be insufficient.
@@ 6.2. Scenario: Detectives
In this scenario, participants continue from a story prefix that describes Detective Wilson’s partner as either trustworthy or untrustworthy. The model then suggests fem-coded or masc-coded language to describe Detective Wilson’s partner (See examples in Table 5). In both cao-etal-2022-theory and our post-survey, participants viewed men as less trustworthy or warm than women (though we note that this difference is not statistically significant in Cao et. al. See Table 16). We therefore treat stories with masc-coded detectives to be pro-stereotypical and stories with trustworthy fem-coded detectives (and untrustworthy masc-coded detectives) to be more pro-stereotypical than their untrustworthy fem-coded (and trustworthy masc-coded) counterparts.
@@ 6.2.1. Effects on Gender Alone
First, we consider the effects of suggestion on gender, regardless of the trustworthiness of Detective Wilson’s partner. Participants specified the partner character’s gender in about $92\%$ of stories (in both treatment and control conditions). We first analyze gender at the story-level, then the relationship between gender and word-level reliance.
At the Story Level
First, we compare the proportion of stories with partners of a given gender that were written with vs without suggestions (Figure 7). Here, we see no significant differences when comparing gender rates without suggestions to rates with masc-coded suggestions (masc-coded partners: $t(242)=0.899$, $p_{\text{FDR}}\approx 0.5681$, $d=0.126$; fem-coded partners: t(242)=1.012t(242)=-1.012, $p_{\text{FDR}}\approx 0.5254$, d=0.141d=-0.141). When the model suggests that the partner should be fem-coded, we see significantly fewer masc-coded partners (t(234)=4.0t(234)=-4.0, $p_{\text{FDR}}\approx 0.0007$, d=0.563d=-0.563) and significantly more fem-coded partners ($t(234)=4.191$, $p_{\text{FDR}}\approx 0.0003$, $d=0.59$). However, even with these changes, we still see significantly more masc-coded partners than fem-coded in all conditions (no suggestions: t(144)=10.934t(144)=-10.934, $p_{\text{FDR}}<0.0001$, d=1.81d=-1.81; masc-coded suggestions: t(340)=21.727t(340)=-21.727, $p_{\text{FDR}}<0.0001$, d=2.35d=-2.35; fem-coded suggestions: t(324)=2.575t(324)=-2.575, $p_{\text{FDR}}\approx 0.0408$, d=0.285d=-0.285).
@@
Bar charts showing the gender distribution of detective partners across no-suggestion, masculine-coded, and feminine-coded suggestion conditions, further grouped by partner trustworthiness. The charts show more masculine-coded than feminine-coded detectives in all conditions, with higher proportions of feminine-coded detectives under feminine-coded suggestions, and higher masculine-coded proportions among untrustworthy partners.
@@
We further assess participants’ reliance on model suggestions at the word level (Table 3). When considering all words suggested by the model in the fem-coded and masc-coded conditions, we see that participants are significantly more likely to write new words or edit model suggestions in the fem-coded detective setting (H2; $t(7091)=4.724$, $p_{\text{FDR}}<0.0001$, $d=0.112$), but we see no significant trend when we constrain this to only the story words that determine the second detective’s inferred gender (H2a; t(302)=0.069t(302)=-0.069, $p_{\text{FDR}}\approx 0.9634$, d=0.008d=-0.008). We also see that participants are significantly less likely to accept model suggestions that would make the second detective fem-coded (H2b; $t(834)=6.729$, $p_{\text{FDR}}<0.0001$, $d=0.479$).